Assume Juno’s paid $368,060 in taxes on taxable income of $1,673,000 last year. This year, the firm paid $401,545 in taxes on taxable income of $1,818,586. Assume the tax rates were the same for both years. What are the marginal and average tax rates for this year? Multiple Choice 22 percent; 21 percent 21 percent; 21 percent 23 percent; 22 percent 22 percent; 22 percent 23 percent; 21 percent
Answer #1

Marginal Tax rate =(Taxes this year-Taxes Last year)/(Taxable Income this year-Taxable Income last year)
=(401545-368060)/(1818586-1673000) =23%
Average tax rate this year =Taxes/Taxable income this year =401545/1818586=22%

Hence 23%;22% is correct option (Option c is correct option)